### 08 Feb 2024 v1

# A Survey on Safe Multi-Modal Learning System

Abstract
With the wide deployment of multimodal learning systems (MMLS) in real-world scenarios, safety concerns have become increasingly prominent. The absence of systematic research into their safety is a significant barrier to progress in this field. To bridge the gap, we present the first taxonomy for MMLS safety, identifying four essential pillars of these concerns. Leveraging this taxonomy, we conduct in-depth reviews for each pillar, highlighting key limitations based on the current state of development. Finally, we pinpoint unique challenges in MMLS safety and provide potential directions for future research.
1 Introduction
Multimodal learning has rapidly evolved in the past few years. Compared to unimodal learning, it could more closely mimic human information processing, where multiple modalities, such as text, image, and audio, are incorporated to understand and interact with the world Baltrušaitis et al. (2018).
Despite significant advancements in related research areas, deploying multimodal learning systems (MMLS) in real-world situations, especially in high-stakes environments like healthcare and autonomous driving, naturally brings about safety challenges. These challenges are notably more intricate and acute than in unimodal environments. Firstly, when implemented in real-world situations, MMLS are likely to face various types of uncommon events and risks. These include natural distribution shifts and deliberately crafted attacks intended to disrupt the systems. The complexity introduced by the multimodal setting may exacerbate these concerns, as it not only involves mismatches within each modality but also introduces domain shifts across modalities McKinzie et al. (2023). Moreover, the rich information spanned by multiple modalities opens doors for cross-modal attacks by malicious entities, further increasing MMLS vulnerability Zhang et al. (2022). The growing risk of privacy is another significant concern. The integration of multiple modalities enhances model efficacy but simultaneously provides attackers with additional information like modality alignment details Rao et al. (2023), exacerbating breaches of privacy. Furthermore, the increasing scales of MMLS further aggravate the challenge of pre-trained models inadvertently memorizing sensitive information from their training data Naseh et al. (2023). MMLS are also at high risk of losing control and generating harmful responses such as discriminatory information or misinformation Schlarmann and Hein (2023); Li et al. (2024). Due to their increasing complexity, especially in large-scale MMLS, monitoring these systems become more challenging. This increased difficulty in oversight makes them more vulnerable to manipulation, resulting in a greater likelihood of performing harmful actions compared to their unimodal counterparts.
Figure 1: Taxonomy for Multimodal Learning Safety.
To advance the development of MMLS, it is essential to thoroughly consider these unique safety challenges and develop appropriate strategies. This survey serves to review recent advances in safe MMLS from four primary pillars: robustness, privacy, monitoring, and controllability. For robustness, we explore two main aspects: the resilience of MMLS against distribution shifts, and their robustness in the face of adversarial perturbations and attacks. Regarding privacy, we first highlight some unique privacy challenges inherent in MMLS, followed by a discussion on various paradigms for preserving privacy in these contexts. For monitoring, our focus is twofold. First, we examine anomaly detection, which is crucial for promptly identifying potential hazards. Second, we investigate the reliability of model outputs to enhance the trustworthiness of MMLS. Controllability is divided into two subcategories. The first, interpretability, aims to study and increase the transparency and understandability of MMLS. The second, fairness, addresses inherent bias in these systems to ensure more equitable and fair responses. We also summarize pertinent benchmarks related to these pillars and conclude the paper with open challenges.
In summary, our main contributions are Taxonomy: We present the first taxonomy for MMLS safety; Review: Following this new taxonomy, we conduct a comprehensive review of existing works, highlighting their limitations; and Open Challenges: We identify some unique challenges related to the safety of MMLS and suggest potential prospects.
2 Robustness
Robustness emerges as a critical research problem when addressing safety concerns in MMLS. In this section, we delineate the problem along two dimensions: robustness against distributional shift, referring to the ability of MMLS to adapt to unexpected natural variations; and adversarial robustness, denoting their resilience to deliberately crafted threats.
2.1 Robustness against Distributional Shift
A standard assumption in ML is consistent distribution, but real-world scenarios often show mismatches between training and testing data. In MMLS settings, these mismatches are more complex, involving shifts within and across modalities. Understanding and enhancing MMLS robustness against distribution shifts is crucial for their safe deployment.
2.1.1 Understanding distributional robustness in MMLS
Most prior research on robust MMLS focuses on vision-language models (VLMs), which are susceptible to distribution shifts in both image and text modalities. Yet they generally show greater robustness than unimodal models Fang et al. (2022). Understanding the reasons behind this robustness gain is vital for advancing MMLS safety. For instance, Fang et al. (2022) identifies diverse image training data as the key to CLIP’s robustness, excluding language supervision and contrastive loss function. However, contrasting views exist Xue et al. (2023), suggesting these latter factors also make contributions. Consensus on these findings is yet to be established. Modality mismatch represents a unique distribution shift in MMLS, which occurs when the data modalities during testing differ from those in training. Earlier studies primarily explored the impact of missing modalities on MMLS’s robustness Ma et al. (2022). More recent research starts considering more generalizable settings McKinzie et al. (2023), investigating the effects of both missing and newly added modalities during test time on MMLS robustness.
2.1.2 Improving MMLS distributional robustness
Current efforts to improve the distributional robustness of MMLS often involve employing data augmentation techniques and implementing robust training strategies.
Data Augmentation. Data augmentation techniques, proven effective in unimodal systems for enhancing model robustness, are naturally extended to MMLS. However, applying unimodal strategies independently to each modality risks losing semantic coherence Liu et al. (2022). Consequently, multimodal techniques preserving modality consistency while improving robustness have emerged. For example, RobustMixGen for VLMs uses a pre-synthesis data separation module to separate data into objects and backgrounds. It maintains cross-modal semantics and boosts model robustness by reducing spurious correlations Kim et al. (2023).
Robust Training. Research in robust training strategies is diverse and spans from improving pre-training strategies to adaptation and fine-tuning methods. For instance, a multi-task optimization algorithm has been developed to automatically identify the optimal data fusion strategies, enhancing models’ robustness against cross-modal distribution shifts Ma et al. (2022). Andonian et al. (2022) presents a novel contrastive learning method for pretraining, combining knowledge distillation and soft alignments to improve robustness. Other studies include devising robust adaptation methods. For instance, Wortsman et al. (2022) shows that fusing weights from zero-shot and fine-tuned CLIPs enhances distributional robustness, with further improvements observed through weight averaging from CLIPs fine-tuned under different hyperparameters.
2.2 Adversarial Robustness
MMLS robustness to adversarial attacks faces unique challenges. Malicious users might use the information from multiple modalities to breach security. Cross-modal adversarial attacks, such as subtle alterations in an image affecting text predictions, further complicate the issue.
2.2.1 Multimodal adversarial attack.
Multimodal adversarial attack aims to trick a target MMLS by subtly altering its multimodal inputs. The objective function for multimodal adversarial attacks can be formulated as:
$\small\mathop{\arg\max}_{\{x_{i}^{adv}\}_{i=1}^{n}}\mathcal{L}(\{x_{i}^{adv}\}_{i=1}^{n},y;\theta)\quad s.t.\,||x_{i}^{adv}-x_{i}||_{p}\leq\epsilon_{i}$ (1)
where $\{x_{i}\}_{i=1}^{n}$ denotes the input composed of $n$ modalities, $\{x_{i}^{adv}\}_{i=1}^{n}$ are the generated adversarial samples, and $\epsilon_{i}$ is the perturbation budget for the $i$-th modality. The attacker optimizes $\mathcal{L}$ to induce model mispredictions within permissible budgets. An effective multimodal adversarial attack needs to properly address the modality gap between multiple input data modalities. Zhang et al. (2022) studies multimodal attacks on VLMs, discovering that launching unimodal attacks independently often fails. They then propose Co-Attack, which coordinates image and text attacks through a sequential strategy Another challenge lies in executing effective, downstream-agnostic attacks, especially within the prevalent ’pretrain-then-finetune’ framework for MMLS, as designing specific attacks for each task a pre-trained encoder might tackle is impractical. Pioneering work by Zhou et al. (2023b) proposes a method for generating broadly applicable adversarial noise, ensuring adversarial samples’ significant deviation from decision boundary in the feature space via a topology-deviation-based generative adversarial network.
2.2.2 Enhance adversarial robustness of multimodal models.
Current works on improving the adversarial robustness of MMLS predominantly center around robust fusion strategy and adversarial training.
Robust Fusion and Alignment. The unique modality alignment process of MMLS, compared to unimodal models, has catalyzed extensive research into enhancing their adversarial robustness. For instance, Yang et al. (2021) improves MMLS single-source adversarial robustness by designing a fusion strategy that detects inconsistencies across modalities and executes feature fusion to counteract the perturbed modality. A constrastive learning framework has been proposed that captures both intra- and inter-modality dynamics, filtering out noise present in various modalities and thereby yielding more robust representations Lin and Hu (2022).
Adversarial Training. Adversarial training enhances robustness effectively, yet its application in MMLS presents challenges, including higher computational demands and scalability issues, compared to unimodal settings. To approach these, Gan et al. (2020) adopts the free adversarial training strategy that obtains parameters’ gradients with almost no extra computational expense. Another challenge lies in improving the generalization of adversarial training procedures within the emerging pretrain-then-finetune paradigm for MMLS. An adversarial pre-training strategy for VLMs is introduced in Gan et al. (2020), focusing on adversarial training in the embedding space of each modality rather than directly perturbing image pixels or text tokens, effectively enhancing model performance in a task-agnostic manner.
2.3 Benchmark
The assessment of model distributional robustness commonly employs metrics such as Effective Robustness, along with Relative Robustness and the MultiModal Impact Score to measure average performance drops during distributional shifts Fang et al. (2022); Qiu et al. (2022). Systematic benchmarks are constantly being proposed as well. For instance, Li et al. (2020) integrates 9 VQA datasets to study four generic robustness in VLMs. Recently, benchmarks for large MMLS have emerged, such as the one introduced in Han et al. (2023). This benchmark assesses GPT-4V’s zero-shot generalization on 13 varied datasets, offering insights into its adaptability to distribution shifts.
2.4 Discussion
Despite the ongoing research, a clear understanding of what underpins the robustness of MMLS remains under-explored, highlighting the need for further studies. Current studies tend to concentrate on VLMs, suggesting the necessity of expanding research focus on a broader spectrum of MMLS. Besides, MMLS vulnerability to sophisticated real-world attacks underscores the importance of enhancing their adversarial robustness. This generally involves devising potent attacks and corresponding defenses, both of which must navigate the complexities of cross-modal information interplay in MMLS. Another challenge is ensuring these defenses can be adapted to the unpredictable nature of real-world attacks.
3 Privacy
3.1 Privacy Issues in MMLS
Significant privacy concerns regarding MMLS have emerged due to their high risk of memorizing and leaking private information Chen et al. (2023c); Naseh et al. (2023). Membership Inference Attacks Ko et al. (2023), for example, could reveal an individual’s health status by checking if their record is in a multimodal disease database. Besides, reconstruction attacks may exploit MMLS’s memorized information to extract private data, such as identifying residents in a geospatial system using modality alignment information between street images and resident details Rao et al. (2023). The recent emerging large MMLS following the pretrain-then-finetune paradigm further intensifies privacy concerns. Each stage of these models, pre-training, fine-tuning, and prompt interaction, carries a high risk of privacy leakage Rao et al. (2023).
3.2 Preserving Privacy in MMLS
Relevant research primarily focuses on differential privacy, machine unlearning, and federated learning.
Differentially private training.
The differential privacy (DP) technique, recognized as a gold standard in privacy preservation, ensures data protection by introducing randomness into its computational processes. DP in MMLS remains understudied. Multimodal DP was first introduced to address the heightened risk of privacy leakage arising from MMLS memorizing sensitive training data information Huang et al. (2023). In this study, DP-CLIP is proposed to safeguard vision-language tasks by learning differentially private image and text representations. It employs per-batch, instead of per-sample, noisy stochastic gradient descent to train the encoder, considering that the CLIP loss function requires contrasting data from different pairs.
Machine unlearning method.
Machine unlearning aims at selectively erasing information (e.g., sensitive information) from a trained model without the need for complete retraining. Progress in machine unlearning has advanced in unimodal learning, but its application in MMLS is still in its infancy. Compared to unimodal scenarios, the key challenge lies in effectively addressing the inherent interdependencies across different data modalities. One work Shaik et al. (2023) presents a modality-specific attention mechanism that generates individual attention scores per modality, averages them across modalities for each data point to determine its final score, and then decides on data point removal and model parameter updates based on this score. Recently a more comprehensive framework for multimodal unlearning has been proposed Cheng and Amiri (2023), encompassing modality decoupling, unimodal knowledge retention, and multimodal knowledge retention. Its first module ensures that individual data modality in deleted pairs are treated as unrelated by the model, while the latter two focus on preserving essential knowledge to maintain the model performance.
Federated learning.
Federated learning enables collaborative model training without the need to share clients’ local data, offering a more privacy-enhancing approach compared to centralized training methods. The MMFed framework, proposed in Xiong et al. (2022), addresses modality discrepancies in MMLS settings through a unified multimodal federated learning approach. It utilizes a cross-attention mechanism to identify key correlations across modalities on local clients, then uploads the parameters of both the classifier and attention module to a central server. CreamFL further advances multimodal federated learning by enabling larger server models and allowing aggregation of clients with diverse architectures and data modalities, thereby aligning more closely with practical scenarios Yu et al. (2022).
3.3 Benchmark
Current measurements of memorization lack standardization and are primarily restricted to VLMs, including metrics like Frechet Inception Distance score Somepalli et al. (2023) and Memorization score Zhang et al. (2023a). Machine unlearning evaluation parallels that in unimodal settings, including attack-based methods like Membership Inference Robustness, and accuracy-based metrics like Mean Squared Error, alongside others like Activation Distance and Retraining Time Nguyen et al. (2022). Benchmarking multimodal federated learning is still in its early stage. FedMultimodal is the first systematic benchmark Feng et al. (2023), offering 10 representative datasets from five key MMLS applications (emotion recognition, multimodal action recognition, human activity recognition, healthcare, and social media) and providing a framework covering the complete pipeline of multimodal federated learning for effective assessment.
3.4 Discussion
Due to the diversity and richness of input data modalities, potential risk of privacy leakage is concerning in MMLS. This involves issues such as sensitive information removed from unimodality may reappear after multimodal fusions. Besides, the increasing scale of MMLS further exacerbates the challenge of memorization. Despite these concerns, privacy issues and their underlying causes are still not thoroughly explored, lacking formal conceptualizations and investigations.
4 Anomaly Detection
Monitoring is a key pillar of AI safety. Here, we discuss an important step for monitoring system behavior: detecting anomalies. While anomaly detection (AD) has been extensively discussed in various fields, MMLS introduce extra complexity owing to heterogeneous modality information, making Multimodal Anomaly Detection (MAD) more challenging. Our review of MAD includes MAD assumptions, research techniques, and benchmarks. We categorize different techniques by the source of modality information.
4.1 Multimodal Anomaly Detection
4.1.1 Vision-Language MAD
In visual modalities, diverse factors like color, light and shooting angles of images can lead to significant variations among samples, resulting in latent space sparsity. This variation affects task performance by disrupting the correlation structure among samples. In contrast, the language modality, characterized by more uniform data, performs well. Existing research predominantly focuses on unimodality, particularly vision, overlooking the richness of multimodal information. Integrating vision and language introduces challenges akin to those faced in unimodal vision, emphasizing the importance of addressing these challenges for improved performance.
[…]
$\small\begin{split}P(f(x),g(s_{i}))=\frac{\exp(<f(x),g(s_{i})>/\tau)}{\sum_{s\in S}\exp(<f(x),g(s_{i})>/\tau)}\end{split}$ (2)
where $\tau>0$ is a hyperparameter. In CLIP, the text prompt template commonly looks like “a photo of a [c]”, where [c] represents the target class name. This differs from vision tasks that involve many objects and use objects’ names as the class name [c]. WinCLIP Jeong et al. (2023) employs anomaly text prompts and calculates the corresponding scores using Eq.2. For prompt optimization, methods such as Dice, and Focal loss are utilized. Subsequent researches mainly follows this approach to compute anomaly scoresZhou et al. (2023a); Chen et al. (2023b).
Adding high-quality anomaly description. Most existing MAD methods only provide anomaly scores and require manual threshold to differentiate between normal and abnormal samples, limiting wider use. Incorporating high-quality anomaly descriptions enhances comprehensive understanding and facilitates practical applications, such as generating training data through simulated anomalous images. AnomalyGPT Gu et al. (2023) designs an MMLS-based image decoder to create an anomaly map and employs prompt embedding for fine-tuning. It supports multi-turn dialogues and exhibits impressive few-shot in-context learning capabilities. However, AnomalyGPT fails to utilize the vision comprehension capacity of MMLS. Myriad Li et al. (2023) addresses this by skillfully incorporating domain knowledge into embeddings, aligning vision-language features, and introducing an Expert Perception module. This module embeds prior knowledge from vision experts as tokens intelligible to LLMs, ensuring precise anomaly detection with detailed descriptions, including the anomaly location, type, and explanation of the judgment.
4.1.2 Video Surveillance MAD
Another common data modality is video. Video Surveillance, also referred to as Closed-circuit television, plays a crucial role in monitoring designated areas for security. Video Surveillance is often treated as an AD problem due to the substantial amount of unlabeled data. Meanwhile, multimodal information in video frames is often underutilized, indicating the need for more effective use. Cutting-edge video MAD methods involve incorporating multimodal signals. For example, MSAF Wei et al. (2022) integrates RGB, optical-flow, and audio modalities, transforming video-level ground truth into pseudo clip-level labels for subsequent training, thereby aligning diverse information to enhance feature learning. Unlike current works that directly feed extracted features into the weakly supervised classifier for frame-level classification, VadCLIP Wu et al. (2023) devises an approach that directly leverages the frozen CLIP without any pre-training or fine-tuning and utilizes multimodal associations between video features and fine-grained language-image alignment information. Video-AnomalyCLIP Zanella et al. (2023) further explore the potential of CLIP, effectively learning text-driven directions for abnormal events by manipulating the latent feature space and projecting anomalous frames onto these directions to exhibit large feature magnitudes for a particular class.
4.2 Benchmark
When evaluating Vision-Language MAD, commonly used benchmark datasets include MVTec-AD, VisA, and MTD. These datasets are annotated at the pixel level and cover a wide range of object subsets. Datasets for Video Surveillance MAD include ShanghaiTech and XD-Violence, an audiovisual abnormal dataset. Performance evaluation is primarily based on metrics such as the area under the ROC curve, area under the precision-recall curve, and F1 score. Other metrics include Per-Region Overlap scores, which assesses anomaly segmentation performance, and Average Precision.
4.3 Discussion
Existing MAD works primarily center around CLIP, resulting in limited exploration of broader multimodal applications. Moreover, current efforts predominantly focus on addressing unsupervised or weakly-supervised MAD challenges, while overlooking the issue of handling incomplete and inaccurate anomaly labels. Notably, limited efforts have been put in studying supervised MAD, where labeling costs might be high. Additionally, while AD has been applied to diverse scenarios like financial fraud and medical systems, the applications of MAD remain limited.
5 Reliable Model Outputs
Ensuring MMLS’s safe deployment requires establishing monitoring mechanisms to assess model trustworthiness. Existing ML methods face challenges in accurately conveying their understanding and providing reliable responses. In MMLS, these issues may become more pronounced given their increased complexity. We discuss this problem from two perspectives: calibration and uncertainty quantification, determining the trust level of MMLS outputs, and red teaming, identifying and mitigating harmful MMLS outputs.
5.1 Calibration and Uncertainty Estimation
Miscalibration in MMLS.
Similar to uni-modal learning, mis-calibration is a prevalent issue in MMLS yet remains underexplored. Despite the involvement of various modalities, MMLS may exhibit a bias towards relying on a single modality for decision-making Ma et al. (2023). Recently, few pilot studies have emerged focusing on foundation MMLS. For example, concerns about miscalibration have been raised regarding zero-shot VLMs such as CLIP. Oh et al. (2023) conducts further studies on fine-tuned VLMs, uncovering that traditional fine-tuning strategies can significantly affect model calibration. Ensuring MMLS are well-calibrated is vital for their safe use, with research concentrating on model calibration and reliable uncertainty estimation (specifically when models are not well-calibrated).
Calibrating MMLS.
The diversity of modalities in MMLS presents both challenges and potential breakthroughs for the development of new approaches for calibrating MMLS. A regularization-based technique CML is proposed for better calibration of MMLS, which calibrates the relationship between confidence and the number of modalities involved in multimodal learning Ma et al. (2023).
Estimating uncertainty in MMLS.
Recent uncertainty estimation research primarily focuses on Bayesian and Conformal Prediction (CP) methods, with extensive work in unimodal settings but limited studies in multimodal contexts. Bayesian-based approaches place a distribution over model parameters and employ marginalization for predictive distribution. Zhang et al. (2023b) pioneers the application of Dempster-Shafer theory to model epistemic uncertainty for each modality and introduces an uncertainty-aware weighting method for dynamic modality fusion. While Bayesian-based approaches are widely used, their computational intensity renders them resource-heavy. Conversely, CP-based methods have emerged as scalable, efficient, and model-agnostic alternatives for uncertainty estimation. Currently, the application of CP in MMLS is limited. Dutta et al. (2023) proposes a CP procedure for foundation models using multimodal web data. It generates plausibility scores based on modality alignment, and then uses them for Monte Carlo-based CP procedure.
5.2 Red Teaming MMLS
The development of foundation models introduces a safety challenge termed ’jailbreak’, where malicious users attempt to circumvent safeguards and instruct the model to produce harmful and unethical responses. Consequently, devising strategies to identify and mitigate such phenomena is vital for safety, a primary focus of red teaming efforts Qi et al. (2023). Historically, research has concentrated on unimodal systems, mostly large language models. There has been a recent shift towards exploring these issues within multimodal contexts. However, research in this area is still nascent, primarily focusing on VLMs. For instance, Qi et al. (2023) designs a simple attack that maximizes the generation probability of harmful responses, effectively jailbreaking VLMs and uncovering MMLS’s vulnerabilities. More recent efforts, such as Li et al. (2024) have begun to establish systematic benchmarks for red teaming in MMLS, paving the way for more in-depth future studies.
5.3 Discussion
The field of ensuring MMLS’s reliable output is crucial yet currently underexplored. Advancing it requires first addressing MMLS’s common miscalibrations, such as overconfidence or underconfidence, then developing sophisticated calibration and uncertainty quantification methods to improve reliability. Besides, red teaming MMLS presents a vital research area. Investigating MMLS’s susceptibility to jailbreak attacks and developing countermeasures against such vulnerabilities offers a promising direction for future research.
6 Interpretability
Being able to comprehend and explain the decisions made by the system is critical for the controllability of MMLS. In this section, we offer a concise overview of Multimodal Explainable Artificial Intelligence (MXAI) Rodis et al. (2023), categorizing MXAI into ante-hoc and post-hoc approaches.
6.1 Multimodal XAI
Ante-hoc MXAI.
Ante-hoc methods integrate interpretability modules directly into the primary model for simultaneous learning. In multimodal contexts, emphasizing the modeling of cross-modal interactions and assessing modalities’ relevance is essential. For example, Li et al. (2018) introduces a multitask learning approach for VQA that concurrently trains an answer predictor and a textual explanation generator, leveraging attention mechanism for cross-modal dynamics. Recently, research on multimodal foundation models has emerged. Wang et al. (2023) devises an interactive multi-agent VQA framework based on VLMs. This framework consists of a seeker and responder that actively generate and exchange information and candidate responses related to the query, and an integrator that synthesizes candidates and generated hypotheses from the seeker and responder to formulate the final answer.
Post-hoc MXAI.
Post-hoc methods generally refers to training an interpretability module based on the output of the primary model. To address fundamental challenges of capturing interactions across modalities, DIME Lyu et al. (2022) proposes to disentangle the black-box model into unimodal contributions and multimodal interactions. This allows for fine-grained analysis of multimodal models while maintaining generality across arbitrary modalities, model architectures, and tasks. Another post-hoc method Panousis and Chatzis (2023) focuses on understanding individual neuron functionality in tasks, utilizing vision-text models and unique network layers to achieve high activation sparsity. It enables the generation of simple textual descriptions, facilitating a direct investigation of the network’s decision process.
6.2 Benchmark
MXAI datasets are typically organized based on tasks, I/O types, dates, and brief descriptions. In evaluating MXAI, the focus is on assessing the effectiveness of explanations in elucidating model decisions. Evaluation methods can be categorized into three types: application-grounded, involving domain experts for real-world utility; human-grounded, testing understandability with individuals regardless of expertise; and functionally-grounded, using formal criteria to assess explanation quality. Detailed descriptions of MXAI benchmarks can be found in Rodis et al. (2023).
6.3 Discussion
Formalizing definitions and terminology is crucial for the MXAI field’s development, as current informal practices and inconsistent definitions hinder progress. Specifically, it is crucial to define explanations precisely and establish standardization for measuring MXAI’s efficiency to enable comparative evaluation of different methods. Another significant challenge arises from the absence of ground-truth modality explanations. For instance, in contrast to textual explanations, the lack of corresponding ground-truth visual explanation data hampers the development of related explanation methods, especially in supervised learning scenarios.
7 Fairness
MMLS incorporates data from various sources like image, text, and audio, presenting unique fairness challenges. Prior work indicates that biases in MMLS stem from individual modalities, which could then be amplified during later modality fusion process Yan et al. (2020). It is crucial to identify and address the specific fairness challenges and underlying causes in MMLS to effectively mitigate discrimination.
7.1 Improving Fairness in MMLS
Addressing fairness in multimodal scenarios is more challenging compared to unimodal due to cross-modality correlation. Given the limited research on MMLS fairness, we categorize fairness algorithms based on whether adjustments occur during model training, grouping pre-processing and post-processing together, separate from in-processing techniques.
7.1.1 Addressing data bias
Pre-processing techniques operate on original input data to remove its inherent bias. For instance, in Automated Video Interviews (AVIs) Booth et al. (2021), researchers tackle gender bias in verbal, paraverbal, and visual communication. They introduces two effective strategies, namely gender-norming and iterative gendered predictor removal, to erase gender-related information from the feature set. Post-processing techniques are performed after training by accessing a holdout set which is not involved during model training. For instance, Chen et al. (2020) determines group-specific random thresholds based on the intersection of group-specific ROC curves. By adding this to the clinical MMLS outputs, the prediction bias can be removed. Notably, post-processing methods are exclusively usable in black-box settings where the model and training data are inaccessible Although pre-processing and post-processing are more generalizable, easier to implement, and less time-consuming, they may lead to uncertain outcomes if not customized for specific algorithms.
7.1.2 Mitigating bias during training
In-processing techniques involve devising training algorithms to mitigate discrimination. Compared to above techniques, in-processing methods tend to achieve superior trade-off between fairness and task performance. For instance, FairCVtest Pena et al. (2020) offers a fictional recruitment testbed demonstrating how biases can be affected by unimodal fators. It further ensures fair multimodal outcomes by generating agnostic representations based on SensitiveNets learning strategy. However, addressing unfairness in multimodal modeling involves more than just tackling unimodal biases. FMMRec Chen et al. (2023a) finds that multimodal recommend systems’ bias stems from repetitively identifying sensitive information across modalities. Adversarial learning method is then employed to disentangle and filter out sensitive information from modality representations. Other approaches involve integrating the aforementioned techniques. For instance, Yan et al. (2020) mitigates gender and race biases through a combination of adversarial learning and data balancing, employing both in-processing and post-processing strategies.
7.2 Benchmark
Datasets for fair MMLS are limited due to privacy concerns and the inherent complexity of multimodal data. Various fairness notions (e.g., Equalized Odds and Statistical Parity Mehrabi et al. (2021)) in MMLS draw inspiration from concepts established in unimodal settings. Key fairness categories encompass individual fairness, focusing on individual samples, and group fairness, targeting protected groups such as Black or Female Chen et al. (2020). In the context of multilingual multimodal models, additional fairness notions like Multilingual Individual and Multilingual Group Fairness have been introduced Wang et al. (2021). The former necessitates similar semantics across languages, linking text snippets to images, while the latter measures equalized predictive performance across languages.
7.3 Discussions
In contrast to the abundance of datasets available for fairness studies in unimodal scenarios, there is a notable absence of reliable datasets for fairness studies in MMLS. This gap is attributed to the intricate nature of modal information, privacy concerns, and the diverse system considerations related to fairness. Existing work, such as Schmitz et al. (2022), compares biases introduced by different modalities and their combinations, yet falls short of delving into the fundamental causes of multimodal unfairness and proposing effective solutions. Furthermore, there is a crucial need for the introduction of new definitions of fairness. While some notable work has suggested some Wang et al. (2021), it lacks a thorough explanation of why this new definition is necessary or when it should be applied.
8 Open Challenges
Our review highlights several key challenges in MMLS safety as briefly illustrated below. We aim to offer insights that can serve as valuable guidance for future research.
Memorization Problem in MMLS. Memorization in large-scale models is a significant contributor to privacy leakage. While this issue has been extensively explored in unimodal systems, research on memorization within MMLS remains limited. There is an urgent need for a formal definition of this problem and the establishment of unified metrics to measure it, especially given the recent rapid advancement of multimodal foundation models.
Differentially Private Training. Differential privacy techniques demonstrate their effectiveness in safeguarding privacy in unimodal systems. Yet, adapting these techniques for MMLS faces challenges and is in its infancy. Some strategies to enhance differential privacy are to apply modality-specific noise addition and coordinate across modalities by considering the interdependence between modalities and developing mechanisms that preserve privacy across the multimodal model without sacrificing utility.
Calibration and Uncertainty Quantification. Miscalibration in multimodal learning, a critical yet under-explored issue, poses challenges to MMLS safety. Future research should focus on understanding this phenomenon and developing advanced calibration methods. Besides, conformal prediction, known for its scalability and minimal assumptions, emerges as a promising solution for quantifying uncertainty. Developing CP methods tailored explicitly for MMLS could significantly benefit practical applications.
Advancing Datasets for MMLS Safety. The challenges associated with multimodal datasets encompass various aspects, such as the lack of available open-source large models capable of effectively handling multimodal inputs. Current models produce content that lacks interpretability, impeding understanding and trust. Ensuring the quality control of generated data poses difficulties, demanding significant resources for collection efforts. Research in MMLS tends to be limited to specific use cases with an uneven distribution of datasets complicating the generalization of models, particularly in terms of fairness. This mirrors the challenges faced by unimodal models. To address these issues, further research and refinement of public datasets are essential.
Safety Challenges in Multimodal LLMs. Multimodal LLMs (e.g., GPT 4V) are designed to understand and generate content across different modalities. Safeguarding multimodal LLMs is very challenging due to various reasons including the additional safety rift induced by different data modalities, often black-box API access, unknown data sources, and emerging LLM safety issues such as hallucination, data leakage, and jailbreak. However, existing efforts to achieve safe multimodal LLMs have been scarce.

### 22 Mar 2024 v2

@@
-A Survey on Safe Multi-Modal Learning System
-Abstract
-With the wide deployment of multimodal learning systems (MMLS) in real-world scenarios, safety concerns have become increasingly prominent.
-The absence of systematic research into their safety is a significant barrier to progress in this field.
-To bridge the gap, we present the first taxonomy for MMLS safety, identifying four essential pillars of these concerns.
-Leveraging this taxonomy, we conduct in-depth reviews for each pillar, highlighting key limitations based on the current state of development.
-Finally, we pinpoint unique challenges in MMLS safety and provide potential directions for future research.
-1 Introduction
-Multimodal learning has rapidly evolved in the past few years.
-Compared to unimodal learning, it could more closely mimic human information processing, where multiple modalities, such as text, image, and audio, are incorporated to understand and interact with the world Baltrušaitis et al.
-(2018).
-Despite significant advancements in related research areas, deploying multimodal learning systems (MMLS) in real-world situations, especially in high-stakes environments like healthcare and autonomous driving, naturally brings about safety challenges.
-These challenges are notably more intricate and acute than in unimodal environments.
+A Survey on Safe Multi-Modal Learning Systems
+Abstract.
+In the rapidly evolving landscape of artificial intelligence, multimodal learning systems (MMLS) have gained traction for their ability to process and integrate information from diverse modality inputs.
+Their expanding use in vital sectors such as healthcare has made safety assurance a critical concern.
+However, the absence of systematic research into their safety is a significant barrier to progress in this field.
+To bridge the gap, we present the first taxonomy that systematically categorizes and assesses MMLS safety.
+This taxonomy is structured around four fundamental pillars that are critical to ensuring the safety of MMLS: robustness, alignment, monitoring, and controllability.
+Leveraging this taxonomy, we review existing methodologies, benchmarks, and the current state of research, while also pinpointing the principal limitations and gaps in knowledge.
+Finally, we discuss unique challenges in MMLS safety.
+In illuminating these challenges, we aim to pave the way for future research, proposing potential directions that could lead to significant advancements in the safety protocols of MMLS.
+Multi-modal System, AI Safety, Responsible AI
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+1. Introduction
+Multimodal learning has rapidly evolved across various domains in the past few years.
+Compared to unimodal learning, it could more closely mimic human information processing, where multiple modalities, such as text, image, and audio, are incorporated to understand and interact with the world (Baltrušaitis et al., 2018).
+Despite significant advancements in related research areas, deploying multimodal learning systems (MMLS) in real-world situations, especially in high-stakes environments like healthcare and autonomous driving, naturally brings safety challenges.
+These challenges are notably more intricate and acute than in unimodal environments (Han et al., 2022).
 Firstly, when implemented in real-world situations, MMLS are likely to face various types of uncommon events and risks.
 These include natural distribution shifts and deliberately crafted attacks intended to disrupt the systems.
-The complexity introduced by the multimodal setting may exacerbate these concerns, as it not only involves mismatches within each modality but also introduces domain shifts across modalities McKinzie et al.
-(2023).
-Moreover, the rich information spanned by multiple modalities opens doors for cross-modal attacks by malicious entities, further increasing MMLS vulnerability Zhang et al.
-(2022).
+The complexity introduced by the multimodal setting may exacerbate these concerns, as it not only involves mismatches within each modality but also introduces domain shifts across modalities (McKinzie et al., 2023).
+Moreover, the rich information spanned by multiple modalities opens doors for cross-modal attacks by malicious entities, further increasing MMLS vulnerability (Zhang et al., 2022).
 The growing risk of privacy is another significant concern.
-The integration of multiple modalities enhances model efficacy but simultaneously provides attackers with additional information like modality alignment details Rao et al.
-(2023), exacerbating breaches of privacy.
-Furthermore, the increasing scales of MMLS further aggravate the challenge of pre-trained models inadvertently memorizing sensitive information from their training data Naseh et al.
-(2023).
-MMLS are also at high risk of losing control and generating harmful responses such as discriminatory information or misinformation Schlarmann and Hein (2023); Li et al.
-(2024).
-Due to their increasing complexity, especially in large-scale MMLS, monitoring these systems become more challenging.
+The integration of multiple modalities enhances model efficacy but simultaneously provides attackers with additional information like modality alignment details (Rao et al., 2023), exacerbating breaches of privacy.
+Furthermore, the increasing scales of MMLS further aggravate the challenge of pre-trained models inadvertently memorizing sensitive information from their training data (Naseh et al., 2023).
+MMLS are also at high risk of losing control and generating harmful responses such as discriminatory information or misinformation (Schlarmann and Hein, 2023; Li et al., 2024).
+Due to their increasing complexity, especially in large-scale MMLS, monitoring these systems becomes more challenging.
 This increased difficulty in oversight makes them more vulnerable to manipulation, resulting in a greater likelihood of performing harmful actions compared to their unimodal counterparts.
-Figure 1: Taxonomy for Multimodal Learning Safety.
+Figure 1. Taxonomy for Safety of Multimodal Learning Systems.
 To advance the development of MMLS, it is essential to thoroughly consider these unique safety challenges and develop appropriate strategies.
-This survey serves to review recent advances in safe MMLS from four primary pillars: robustness, privacy, monitoring, and controllability.
+This survey serves to review recent advances in safe MMLS from four primary pillars: robustness, alignment, monitoring, and controllability.
 For robustness, we explore two main aspects: the resilience of MMLS against distribution shifts, and their robustness in the face of adversarial perturbations and attacks.
-Regarding privacy, we first highlight some unique privacy challenges inherent in MMLS, followed by a discussion on various paradigms for preserving privacy in these contexts.
+Regarding alignment, we first examine the issue of the misalignment phenomenon in MMLS, noting their susceptibility to being jailbroken and producing harmful content.
+We then proceed to introduce the current techniques developed for aligning MMLS.
 For monitoring, our focus is twofold.
@@
 Second, we investigate the reliability of model outputs to enhance the trustworthiness of MMLS.
-Controllability is divided into two subcategories.
+Controllability is divided into three subcategories.
 The first, interpretability, aims to study and increase the transparency and understandability of MMLS.
-The second, fairness, addresses inherent bias in these systems to ensure more equitable and fair responses.
-We also summarize pertinent benchmarks related to these pillars and conclude the paper with open challenges.
-In summary, our main contributions are Taxonomy: We present the first taxonomy for MMLS safety; Review: Following this new taxonomy, we conduct a comprehensive review of existing works, highlighting their limitations; and Open Challenges: We identify some unique challenges related to the safety of MMLS and suggest potential prospects.
-2 Robustness
+The second, privacy, discusses unique privacy challenges within MMLS, and various paradigms for preserving privacy in these contexts.
+Lastly, fairness addresses inherent bias in these systems to ensure more equitable and fair responses.
+We also discuss the limitations and summarize pertinent benchmarks related to these pillars and conclude the paper with open challenges.
+In summary, our main contributions are Taxonomy: We present the first taxonomy (Figure 1) for MMLS safety; Review: Following this new taxonomy, we conduct a comprehensive review of existing works, highlighting their limitations; and Open Challenges: We identify some unique challenges related to the safety of MMLS and suggest potential prospects.
+2. Robustness
 Robustness emerges as a critical research problem when addressing safety concerns in MMLS.
 In this section, we delineate the problem along two dimensions: robustness against distributional shift, referring to the ability of MMLS to adapt to unexpected natural variations; and adversarial robustness, denoting their resilience to deliberately crafted threats.
-2.1 Robustness against Distributional Shift
+2.1. Robustness against Distributional Shift
 A standard assumption in ML is consistent distribution, but real-world scenarios often show mismatches between training and testing data.
@@
 Understanding and enhancing MMLS robustness against distribution shifts is crucial for their safe deployment.
-2.1.1 Understanding distributional robustness in MMLS
+2.1.1. Understanding distributional robustness in MMLS
 Most prior research on robust MMLS focuses on vision-language models (VLMs), which are susceptible to distribution shifts in both image and text modalities.
-Yet they generally show greater robustness than unimodal models Fang et al.
-(2022).
+Yet they generally show greater robustness than unimodal models (Fang et al., 2022).
 Understanding the reasons behind this robustness gain is vital for advancing MMLS safety.
-For instance, Fang et al.
-(2022) identifies diverse image training data as the key to CLIP’s robustness, excluding language supervision and contrastive loss function.
-However, contrasting views exist Xue et al.
-(2023), suggesting these latter factors also make contributions.
+For instance, (Fang et al., 2022) identifies diverse image training data as the key to CLIP’s robustness, excluding language supervision and contrastive loss function.
+However, contrasting views exist (Xue et al., 2023), suggesting these latter factors also make contributions.
 Consensus on these findings is yet to be established.
 Modality mismatch represents a unique distribution shift in MMLS, which occurs when the data modalities during testing differ from those in training.
-Earlier studies primarily explored the impact of missing modalities on MMLS’s robustness Ma et al.
-(2022).
-More recent research starts considering more generalizable settings McKinzie et al.
-(2023), investigating the effects of both missing and newly added modalities during test time on MMLS robustness.
-2.1.2 Improving MMLS distributional robustness
-Current efforts to improve the distributional robustness of MMLS often involve employing data augmentation techniques and implementing robust training strategies.
+Earlier studies primarily explored the impact of missing modalities on MMLS’s robustness (Ma et al., 2022).
+More recent research starts considering more generalizable settings (McKinzie et al., 2023), investigating the effects of both missing and newly added modalities during test time on MMLS robustness.
+2.1.2. Improving MMLS distributional robustness
+Current efforts to improve the distributional robustness of MMLS often involve data augmentation techniques and robust training strategies.
 Data Augmentation.
 Data augmentation techniques, proven effective in unimodal systems for enhancing model robustness, are naturally extended to MMLS.
-However, applying unimodal strategies independently to each modality risks losing semantic coherence Liu et al.
-(2022).
+However, applying unimodal strategies independently to each modality risks losing semantic coherence (Liu et al., 2022).
 Consequently, multimodal techniques preserving modality consistency while improving robustness have emerged.
 For example, RobustMixGen for VLMs uses a pre-synthesis data separation module to separate data into objects and backgrounds.
-It maintains cross-modal semantics and boosts model robustness by reducing spurious correlations Kim et al.
-(2023).
+It maintains cross-modal semantics and boosts model robustness by reducing spurious correlations (Kim et al., 2023).
 Robust Training.
 Research in robust training strategies is diverse and spans from improving pre-training strategies to adaptation and fine-tuning methods.
-For instance, a multi-task optimization algorithm has been developed to automatically identify the optimal data fusion strategies, enhancing models’ robustness against cross-modal distribution shifts Ma et al.
-(2022).
-Andonian et al.
-(2022) presents a novel contrastive learning method for pretraining, combining knowledge distillation and soft alignments to improve robustness.
+For instance, a multi-task optimization algorithm has been developed to automatically identify the optimal data fusion strategies, enhancing models’ robustness against cross-modal distribution shifts (Ma et al., 2022).
+Andonian et al.(Andonian et al., 2022) presents a novel contrastive learning method for pretraining, combining knowledge distillation and soft alignments to improve robustness.
 Other studies include devising robust adaptation methods.
-For instance, Wortsman et al.
-(2022) shows that fusing weights from zero-shot and fine-tuned CLIPs enhances distributional robustness, with further improvements observed through weight averaging from CLIPs fine-tuned under different hyperparameters.
-2.2 Adversarial Robustness
+For instance, (Wortsman et al., 2022) shows that fusing weights from zero-shot and fine-tuned CLIPs enhances distributional robustness, with further improvements observed through weight averaging from CLIPs fine-tuned under different hyperparameters.
+2.2. Adversarial Robustness
 MMLS robustness to adversarial attacks faces unique challenges.
 Malicious users might use the information from multiple modalities to breach security.
-Cross-modal adversarial attacks, such as subtle alterations in an image affecting text predictions, further complicate the issue.
-2.2.1 Multimodal adversarial attack.
-Multimodal adversarial attack aims to trick a target MMLS by subtly altering its multimodal inputs. The objective function for multimodal adversarial attacks can be formulated as:
-$\small\mathop{\arg\max}_{\{x_{i}^{adv}\}_{i=1}^{n}}\mathcal{L}(\{x_{i}^{adv}\}_{i=1}^{n},y;\theta)\quad s.t.\,||x_{i}^{adv}-x_{i}||_{p}\leq\epsilon_{i}$ (1)
+Cross-modal adversarial attacks, such as subtle alterations in an image affecting text predictions, further complicate the issue (Zhang et al., 2022).
+2.2.1. Multimodal adversarial attack.
+Multimodal adversarial attack aims to trick a target MMLS by subtly altering its multimodal inputs.
+The objective function for multimodal adversarial attacks can generally be formulated as (Tian and Xu, 2021):
+(1) $\mathop{\arg\max}_{\{x_{i}^{adv}\}_{i=1}^{n}}\mathcal{L}(\{x_{i}^{adv}\}_{i=1}^{n},y;\theta)\quad s.t.\,||x_{i}^{adv}-x_{i}||_{p}\leq\epsilon_{i},$
 where $\{x_{i}\}_{i=1}^{n}$ denotes the input composed of $n$ modalities, $\{x_{i}^{adv}\}_{i=1}^{n}$ are the generated adversarial samples, and $\epsilon_{i}$ is the perturbation budget for the $i$-th modality.
-The attacker optimizes $\mathcal{L}$ to induce model mispredictions within permissible budgets.
+The attacker optimizes $\mathcal{L}$ parameterized by $\theta$ to induce model mispredictions within permissible budgets.
 An effective multimodal adversarial attack needs to properly address the modality gap between multiple input data modalities.
-Zhang et al.
-(2022) studies multimodal attacks on VLMs, discovering that launching unimodal attacks independently often fails.
-They then propose Co-Attack, which coordinates image and text attacks through a sequential strategy Another challenge lies in executing effective, downstream-agnostic attacks, especially within the prevalent ’pretrain-then-finetune’ framework for MMLS, as designing specific attacks for each task a pre-trained encoder might tackle is impractical.
-Pioneering work by Zhou et al.
-(2023b) proposes a method for generating broadly applicable adversarial noise, ensuring adversarial samples’ significant deviation from decision boundary in the feature space via a topology-deviation-based generative adversarial network.
-2.2.2 Enhance adversarial robustness of multimodal models.
-Current works on improving the adversarial robustness of MMLS predominantly center around robust fusion strategy and adversarial training.
+(Zhang et al., 2022) studies multimodal attacks on VLMs, discovering that launching unimodal attacks independently often fails.
+They then propose Co-Attack, which coordinates image and text attacks through a sequential strategy Another challenge lies in executing effective, downstream-agnostic attacks, especially within the prevalent ‘pretrain-then-finetune’ framework for MMLS, as designing specific attacks for a pre-trained encoder targeting each task is impractical.
+Pioneering work by (Zhou et al., 2023a) proposes a method for generating broadly applicable adversarial noise, ensuring adversarial samples’ significant deviation from decision boundary in the feature space via a topology-deviation-based generative adversarial network.
+2.2.2. Enhance adversarial robustness of multimodal learning models.
+Current works on improving the adversarial robustness of MMLS center around robust fusion strategy and adversarial training.
 Robust Fusion and Alignment.
 The unique modality alignment process of MMLS, compared to unimodal models, has catalyzed extensive research into enhancing their adversarial robustness.
-For instance, Yang et al.
-(2021) improves MMLS single-source adversarial robustness by designing a fusion strategy that detects inconsistencies across modalities and executes feature fusion to counteract the perturbed modality.
-A constrastive learning framework has been proposed that captures both intra- and inter-modality dynamics, filtering out noise present in various modalities and thereby yielding more robust representations Lin and Hu (2022).
+For instance, (Yang et al., 2022) improves MMLS single-source adversarial robustness by designing a fusion strategy that detects inconsistencies across modalities and executes feature fusion to counteract the perturbed modality.
+A constrastive learning framework has been proposed that captures both intra- and inter-modality dynamics, filtering out noise present in various modalities and thereby yielding more robust representations (Lin and Hu, 2022).
 Adversarial Training.
-Adversarial training enhances robustness effectively, yet its application in MMLS presents challenges, including higher computational demands and scalability issues, compared to unimodal settings.
-To approach these, Gan et al.
-(2020) adopts the free adversarial training strategy that obtains parameters’ gradients with almost no extra computational expense.
+Adversarial training enhances robustness effectively, yet its application in MMLS presents challenges, including higher computational demands and scalability issues.
+To approach these, (Gan et al., 2020) adopts the free adversarial training strategy that obtains parameters’ gradients with almost no extra computational expense.
 Another challenge lies in improving the generalization of adversarial training procedures within the emerging pretrain-then-finetune paradigm for MMLS.
-An adversarial pre-training strategy for VLMs is introduced in Gan et al.
-(2020), focusing on adversarial training in the embedding space of each modality rather than directly perturbing image pixels or text tokens, effectively enhancing model performance in a task-agnostic manner.
-2.3 Benchmark
-The assessment of model distributional robustness commonly employs metrics such as Effective Robustness, along with Relative Robustness and the MultiModal Impact Score to measure average performance drops during distributional shifts Fang et al.
-(2022); Qiu et al.
-(2022).
+An adversarial pre-training strategy for VLMs is introduced in (Gan et al., 2020), focusing on adversarial training in the embedding space of each modality rather than directly perturbing image pixels or text tokens, effectively enhancing model performance in a task-agnostic manner.
+2.3. Benchmark
+The assessment of model distributional robustness commonly employs metrics such as Effective Robustness, along with Relative Robustness and the MultiModal Impact Score to measure average performance drops during distributional shifts (Fang et al., 2022; Qiu et al., 2022).
 Systematic benchmarks are constantly being proposed as well.
-For instance, Li et al.
-(2020) integrates 9 VQA datasets to study four generic robustness in VLMs.
-Recently, benchmarks for large MMLS have emerged, such as the one introduced in Han et al.
-(2023).
-This benchmark assesses GPT-4V’s zero-shot generalization on 13 varied datasets, offering insights into its adaptability to distribution shifts.
-2.4 Discussion
+For instance, (Li et al., 2020) integrates nine VQA datasets to study four generic robustness in VLMs.
+Recently, benchmarks for large MMLS have emerged, such as the one introduced in (Han et al., 2023).
+This benchmark assesses GPT-4V’s (Achiam et al., 2023) zero-shot generalization on 13 varied datasets, offering insights into its adaptability to distribution shifts.
+2.4. Discussion
 Despite the ongoing research, a clear understanding of what underpins the robustness of MMLS remains under-explored, highlighting the need for further studies.
@@
 Another challenge is ensuring these defenses can be adapted to the unpredictable nature of real-world attacks.
-3 Privacy
-3.1 Privacy Issues in MMLS
-Significant privacy concerns regarding MMLS have emerged due to their high risk of memorizing and leaking private information Chen et al.
-(2023c); Naseh et al.
-(2023).
-Membership Inference Attacks Ko et al.
-(2023), for example, could reveal an individual’s health status by checking if their record is in a multimodal disease database.
-Besides, reconstruction attacks may exploit MMLS’s memorized information to extract private data, such as identifying residents in a geospatial system using modality alignment information between street images and resident details Rao et al.
-(2023).
-The recent emerging large MMLS following the pretrain-then-finetune paradigm further intensifies privacy concerns.
-Each stage of these models, pre-training, fine-tuning, and prompt interaction, carries a high risk of privacy leakage Rao et al.
-(2023).
-3.2 Preserving Privacy in MMLS
-Relevant research primarily focuses on differential privacy, machine unlearning, and federated learning.
-Differentially private training.
-The differential privacy (DP) technique, recognized as a gold standard in privacy preservation, ensures data protection by introducing randomness into its computational processes.
-DP in MMLS remains understudied.
-Multimodal DP was first introduced to address the heightened risk of privacy leakage arising from MMLS memorizing sensitive training data information Huang et al.
-(2023).
-In this study, DP-CLIP is proposed to safeguard vision-language tasks by learning differentially private image and text representations.
-It employs per-batch, instead of per-sample, noisy stochastic gradient descent to train the encoder, considering that the CLIP loss function requires contrasting data from different pairs.
-Machine unlearning method.
-Machine unlearning aims at selectively erasing information (e.g., sensitive information) from a trained model without the need for complete retraining.
-Progress in machine unlearning has advanced in unimodal learning, but its application in MMLS is still in its infancy.
-Compared to unimodal scenarios, the key challenge lies in effectively addressing the inherent interdependencies across different data modalities.
-One work Shaik et al.
-(2023) presents a modality-specific attention mechanism that generates individual attention scores per modality, averages them across modalities for each data point to determine its final score, and then decides on data point removal and model parameter updates based on this score.
-Recently a more comprehensive framework for multimodal unlearning has been proposed Cheng and Amiri (2023), encompassing modality decoupling, unimodal knowledge retention, and multimodal knowledge retention.
-Its first module ensures that individual data modality in deleted pairs are treated as unrelated by the model, while the latter two focus on preserving essential knowledge to maintain the model performance.
-Federated learning.
-Federated learning enables collaborative model training without the need to share clients’ local data, offering a more privacy-enhancing approach compared to centralized training methods.
-The MMFed framework, proposed in Xiong et al.
-(2022), addresses modality discrepancies in MMLS settings through a unified multimodal federated learning approach.
-It utilizes a cross-attention mechanism to identify key correlations across modalities on local clients, then uploads the parameters of both the classifier and attention module to a central server.
-CreamFL further advances multimodal federated learning by enabling larger server models and allowing aggregation of clients with diverse architectures and data modalities, thereby aligning more closely with practical scenarios Yu et al.
-(2022).
-3.3 Benchmark
-Current measurements of memorization lack standardization and are primarily restricted to VLMs, including metrics like Frechet Inception Distance score Somepalli et al.
-(2023) and Memorization score Zhang et al.
-(2023a).
-Machine unlearning evaluation parallels that in unimodal settings, including attack-based methods like Membership Inference Robustness, and accuracy-based metrics like Mean Squared Error, alongside others like Activation Distance and Retraining Time Nguyen et al.
-(2022).
-Benchmarking multimodal federated learning is still in its early stage.
-FedMultimodal is the first systematic benchmark Feng et al.
-(2023), offering 10 representative datasets from five key MMLS applications (emotion recognition, multimodal action recognition, human activity recognition, healthcare, and social media) and providing a framework covering the complete pipeline of multimodal federated learning for effective assessment.
-3.4 Discussion
-Due to the diversity and richness of input data modalities, potential risk of privacy leakage is concerning in MMLS.
-This involves issues such as sensitive information removed from unimodality may reappear after multimodal fusions.
-Besides, the increasing scale of MMLS further exacerbates the challenge of memorization.
-Despite these concerns, privacy issues and their underlying causes are still not thoroughly explored, lacking formal conceptualizations and investigations.
-4 Anomaly Detection
-Monitoring is a key pillar of AI safety.
+3. Alignment
+Aligning MMLS with human values is a crucial yet challenging aspect of safe deployment.
+Recent studies have highlighted vulnerabilities in current foundation models, revealing that they can be manipulated to yield harmful and misaligned outputs.
+This issue could be further exacerbated in MMLS than unimodal foundation models such as LLMs.
+Therefore, delving into this phenomenon and developing more safely-aligned MMLS is a critical area of interest.
+We discuss this issue from two perspectives: the misalignment in MMLS and strategies for their alignment.
+3.1. Misalignment: Jailbreaking MMLS
+The development of foundation models introduces a safety challenge termed ‘jailbreak’, where malicious users attempt to circumvent safeguards and instruct the model to produce harmful responses that are misaligned with ethical standards.
+Previous research has mainly focused on unimodal foundation models, particularly LLMs, with studies on MMLS being relatively new and concentrating largely on VLMs.
+One of the most prevalent jailbreak approaches is adversarial prompting.
+Compared to LLMs, MMLS can be more vulnerable to jailbreak attacks owing to the integration of additional modalities.
+For instance, (Carlini et al., 2024) successfully jailbreaks VLMs via constructing adversarial image input.
+Besides, (Qi et al., 2023) designs a simple attack that maximizes the generation probability of harmful responses, effectively jailbreaking VLMs with a single visual adversarial example.
+More recent work (Niu et al., 2024) investigates the potential of constructing adversarial textual input based on image jailbreak prompt, introducing new approaches to jailbreak MMLS.
+3.2. Aligning MMLS
+At present, the two most common techniques for achieving model alignment are Instruction Tuning (Liu et al., 2024), which finetunes models using collections of pairs of instructions and faithful outputs, and Reinforcement Learning from Human Feedback (RLHF), which employs reinforcement learning techniques to optimize model outputs (Stiennon et al., 2020; Ouyang et al., 2022).
+When aligning MMLS, a significant hurdle is the scarcity of available data.
+Multimodal data pairs are more challenging to gather from the internet compared to textual data alone.
+Another issue is the quality of the data collected.
+For example, an image paired with a brief description is not as informative as one accompanied by a comprehensive narrative.
+LLaVA-RLHF (Sun et al., 2023) marks a pioneering effort in adapting RLHF from purely textual applications to vision-language alignment tasks, using human annotated preferences to fine-tune multimodal models via reinforcement learning.
+It effectively addresses multimodal misalignment at a relatively low cost for annotations.
+(Li et al., 2023b) first constructs a high-quality instruction set based on various multimodal instruction tuning sources and AI annotations, and then employs direct preference optimization (Rafailov et al., 2024) to enhance VLMs’ alignment, resulting in more faithful systems and mitigating misalignment behaviors.
+3.3. Benchmark
+Common evaluation metrics include Attack Success Rate, Recognition Success Rate, and Defense Success Rate (Wu et al., 2023a).
+Recently, researchers have begun to establish systematic benchmarks for measuring misalignment in MMLS.
+For instance, (Li et al., 2024) proposes a multimodal dataset RTVLM for a comprehensive assessment of VLMs across various red-teaming cases.
+(Niu et al., 2024) expands upon AdvBench to construct a multimodal benchmark, AdvBench-M, which consists of eight distinct categories of harmful behaviors for evaluation.
+3.4. Discussion
+Aligning multimodal foundation models is notably more complex because of the involvement of multiple modalities (Qi et al., 2023).
+This complexity increases the potential for vulnerabilities and intensifies the challenge of establishing robust defenses.
+For instance, incorporating visual data into LLMs creates opportunities for adversarial attacks, which can be executed more readily than with purely discrete textual inputs.
+Consequently, this facilitates the jailbreaking of MMLS, often leading to the generation of toxic content, misinformation, and hallucination.
+Although researchers have devised various strategies for aligning MMLS, the effectiveness of these methods is still under scrutiny.
+Even with attempts at safety alignment, MMLS can be compromised with relative ease.
+Thus, the pursuit of alignment within MMLS is fraught with difficulties and remains an area in need of more thorough exploration and understanding.
+4. Monitoring-Anomaly Detection
+Monitoring is crucial for AI safety.
 Here, we discuss an important step for monitoring system behavior: detecting anomalies.
-While anomaly detection (AD) has been extensively discussed in various fields, MMLS introduce extra complexity owing to heterogeneous modality information, making Multimodal Anomaly Detection (MAD) more challenging.
+Anomaly detection (AD) refers to the problem of finding patterns in data that do not conform to expected behavior.
+While AD has been extensively discussed in various fields, MMLS introduce extra complexity owing to heterogeneous modality information, making Multimodal Anomaly Detection (MAD) more challenging.
 Our review of MAD includes MAD assumptions, research techniques, and benchmarks.
-We categorize different techniques by the source of modality information.
-4.1 Multimodal Anomaly Detection
-4.1.1 Vision-Language MAD
-In visual modalities, diverse factors like color, light and shooting angles of images can lead to significant variations among samples, resulting in latent space sparsity.
-This variation affects task performance by disrupting the correlation structure among samples.
-In contrast, the language modality, characterized by more uniform data, performs well.
-Existing research predominantly focuses on unimodality, particularly vision, overlooking the richness of multimodal information.
-Integrating vision and language introduces challenges akin to those faced in unimodal vision, emphasizing the importance of addressing these challenges for improved performance.
+4.1. Failure Detection in MAD
+A common practice of MAD in MMLS is to detect failure, which can serve as an early warning system for potential failures by identifying irregularities.
+Deployment failures of MMLS are prevalent despite rigorous testing before development.
+One of the challenges lies in anticipating and testing all potential failures in advance.
+To bridge this gap, MULTIMON (Tong et al., 2024) is introduced to automatically detect systematic failures in deployed MMLS.
+It first identifies individual failures based on CLIP embeddings, which encode multimodal information, and erroneous agreement, and then leverages language models to describe patterns of failures.
+The framework proposed by (Jain et al., 2022) aims to detect failure modes within the latent space.
+It seeks to discover a hyperplane that optimally distinguishes between correct and incorrect examples, interprets the patterns of errors, and then enhances model reliability during training along the axes of failure.
+Similarly, DOMINO (Eyuboglu et al., 2022) tackles the challenge of identifying underperforming subsets within MMLS, particularly dealing with audiovisual inputs.
+Recent studies (d’Eon et al., 2022; Yeh et al., 2020) have presented Slice Discovery Methods (SDMs) that use model representations to detect underperforming segments, however, these methods often lack a robust framework for quantitative analysis.
+Thus DOMINO has been developed as an extensive evaluation framework for SDMs in various contexts and introduced as an innovative SDM that employs cross-modal embeddings and an error-aware mixture model for improved performance.
+It outperforms prior methods, demonstrating improved coherence and performance in slice discovery.
+4.2. Other Applications of MAD
+4.2.1. Vision-Language MAD
+Vision modality faces sample variations due to factors like color and light, causing latent space sparsity and disrupted correlations, which impact task performance.
+Conversely, the more uniform language modality often performs better.
+Existing research (Pang et al., 2021) mostly focuses on unimodality, particularly vision, overlooking multimodal benefits.
+Meanwhile integrating multimodalities also introuces challenges to be addressed.
 […]
-$\small\begin{split}P(f(x),g(s_{i}))=\frac{\exp(<f(x),g(s_{i})>/\tau)}{\sum_{s\in S}\exp(<f(x),g(s_{i})>/\tau)}\end{split}$ (2)
+(2) $\begin{split}P(f(x),g(s_{i}))=\frac{\exp(<f(x),g(s_{i})>/\tau)}{\sum_{s\in S}\exp(<f(x),g(s_{i})>/\tau)},\end{split}$
 where $\tau>0$ is a hyperparameter.
 In CLIP, the text prompt template commonly looks like “a photo of a [c]”, where [c] represents the target class name.
-This differs from vision tasks that involve many objects and use objects’ names as the class name [c].
-WinCLIP Jeong et al.
-(2023) employs anomaly text prompts and calculates the corresponding scores using Eq.2.
-For prompt optimization, methods such as Dice, and Focal loss are utilized.
-Subsequent researches mainly follows this approach to compute anomaly scoresZhou et al.
-(2023a); Chen et al.
-(2023b).
+It differs from vision tasks that involve many objects and use objects’ names as [c].
+WinCLIP (Jeong et al., 2023) employs anomaly text prompts and calculates the corresponding scores using Eq.2.
+For prompt optimization, a linear combination of the focal loss (Lin et al., 2017) and dice loss (Milletari et al., 2016) is usually used.
+Subsequent research mainly follows this approach to compute anomaly scores (Zhou et al., 2023b; Deng et al., 2023; Chen et al., 2023c).
 Adding high-quality anomaly description.
-Most existing MAD methods only provide anomaly scores and require manual threshold to differentiate between normal and abnormal samples, limiting wider use.
-Incorporating high-quality anomaly descriptions enhances comprehensive understanding and facilitates practical applications, such as generating training data through simulated anomalous images.
-AnomalyGPT Gu et al.
-(2023) designs an MMLS-based image decoder to create an anomaly map and employs prompt embedding for fine-tuning.
+Most existing MAD methods only provide anomaly scores and require manual thresholds to differentiate between normal and abnormal samples, limiting wider use.
+Incorporating high-quality anomaly descriptions can enhance understanding and facilitate practical applications.
+AnomalyGPT (Gu et al., 2023) designs an MMLS-based image decoder for anomaly mapping and employs prompt embedding for fine-tuning.
 It supports multi-turn dialogues and exhibits impressive few-shot in-context learning capabilities.
-However, AnomalyGPT fails to utilize the vision comprehension capacity of MMLS.
-Myriad Li et al.
-(2023) addresses this by skillfully incorporating domain knowledge into embeddings, aligning vision-language features, and introducing an Expert Perception module.
-This module embeds prior knowledge from vision experts as tokens intelligible to LLMs, ensuring precise anomaly detection with detailed descriptions, including the anomaly location, type, and explanation of the judgment.
-4.1.2 Video Surveillance MAD
-Another common data modality is video.
-Video Surveillance, also referred to as Closed-circuit television, plays a crucial role in monitoring designated areas for security.
-Video Surveillance is often treated as an AD problem due to the substantial amount of unlabeled data.
+Yet, it underutilizes MMLS’s vision comprehension potential.
+Myriad (Li et al., 2023a) addresses this by skillfully incorporating domain knowledge into embeddings, aligning vision-language features, and introducing an Expert Perception module.
+This module embeds prior knowledge from vision experts as tokens intelligible to LLMs, ensuring precise AD with detailed descriptions, including the anomaly location, type, and explanation of the judgment.
+4.2.2. Video Surveillance MAD
+Another common multimodal data is video.
+Video Surveillance (Pang et al., 2021) plays a crucial role in monitoring areas for security.
+It is often treated as an AD problem due to the substantial amount of unlabeled data.
 Meanwhile, multimodal information in video frames is often underutilized, indicating the need for more effective use.
 Cutting-edge video MAD methods involve incorporating multimodal signals.
-For example, MSAF Wei et al.
-(2022) integrates RGB, optical-flow, and audio modalities, transforming video-level ground truth into pseudo clip-level labels for subsequent training, thereby aligning diverse information to enhance feature learning.
-Unlike current works that directly feed extracted features into the weakly supervised classifier for frame-level classification, VadCLIP Wu et al.
-(2023) devises an approach that directly leverages the frozen CLIP without any pre-training or fine-tuning and utilizes multimodal associations between video features and fine-grained language-image alignment information.
-Video-AnomalyCLIP Zanella et al.
-(2023) further explore the potential of CLIP, effectively learning text-driven directions for abnormal events by manipulating the latent feature space and projecting anomalous frames onto these directions to exhibit large feature magnitudes for a particular class.
-4.2 Benchmark
-When evaluating Vision-Language MAD, commonly used benchmark datasets include MVTec-AD, VisA, and MTD.
-These datasets are annotated at the pixel level and cover a wide range of object subsets.
-Datasets for Video Surveillance MAD include ShanghaiTech and XD-Violence, an audiovisual abnormal dataset.
+For example, MSAF (Wei et al., 2022) integrates RGB, optical-flow, and audio modalities, transforming video-level ground truth into pseudo clip-level labels for subsequent training, thereby aligning diverse information to enhance learning.
+Unlike current works that directly use extracted features for frame-level classification, VadCLIP (Wu et al., 2023b) devises an approach that leverages the frozen CLIP and multimodal associations between video features and fine-grained language-image alignment information.
+Video-AnomalyCLIP (Zanella et al., 2023) further explore the potential of CLIP, effectively learning text-driven directions for abnormal events by manipulating the latent feature space and projecting anomalous frames onto these directions to exhibit large feature magnitudes for a particular class.
+4.3. Benchmark
+In the domain of failure detection, datasets typically consist of normal data without explicitly labeled anomaly subsets.
+For failure detection in MAD, the choice of datasets varies, reflecting diverse areas of study encompassing medical imaging to image-text corpora.
+Conversely, in AD scenarios, datasets are specifically annotated to highlight anomalies.
+When evaluating Vision-Language MAD, commonly used benchmark datasets include MVTec-AD (Bergmann et al., 2019), VisA (Zou et al., 2022), KSDD2 (Božič et al., 2021), MPDD (Jezek et al., 2021), MTD (Huang et al., 2020), BTAD (Mishra et al., 2021), etc.
+These industrial inspection datasets are annotated at the pixel level and cover a wide range of object subsets.
+Datasets for Video Surveillance MAD include ShanghaiTech (Liu et al., 2018), UCF-Crime (Sultani et al., 2018) and XD-Violence (Wu et al., 2020), an audiovisual abnormal dataset.
 Performance evaluation is primarily based on metrics such as the area under the ROC curve, area under the precision-recall curve, and F1 score.
-Other metrics include Per-Region Overlap scores, which assesses anomaly segmentation performance, and Average Precision.
-4.3 Discussion
-Existing MAD works primarily center around CLIP, resulting in limited exploration of broader multimodal applications.
+Other metrics include Per-Region Overlap scores, which assess anomaly segmentation performance, and Average Precision.
+4.4. Discussion
+Existing MAD works primarily center around CLIP, resulting in limited exploration of broader multimodal applications (Yoo et al., 2021; Wang et al., 2023a; Ji et al., 2020).
 Moreover, current efforts predominantly focus on addressing unsupervised or weakly-supervised MAD challenges, while overlooking the issue of handling incomplete and inaccurate anomaly labels.
-Notably, limited efforts have been put in studying supervised MAD, where labeling costs might be high.
+Notably, limited efforts have been put into studying supervised MAD, where labeling costs might be high.
 Additionally, while AD has been applied to diverse scenarios like financial fraud and medical systems, the applications of MAD remain limited.
-5 Reliable Model Outputs
-Ensuring MMLS’s safe deployment requires establishing monitoring mechanisms to assess model trustworthiness.
+5. Monitoring-Reliable Model Outputs
+Ensuring MMLS’s safe deployment requires establishing monitoring mechanisms to assess model reliability.
 Existing ML methods face challenges in accurately conveying their understanding and providing reliable responses.
 In MMLS, these issues may become more pronounced given their increased complexity.
-We discuss this problem from two perspectives: calibration and uncertainty quantification, determining the trust level of MMLS outputs, and red teaming, identifying and mitigating harmful MMLS outputs.
-5.1 Calibration and Uncertainty Estimation
-Miscalibration in MMLS.
-Similar to uni-modal learning, mis-calibration is a prevalent issue in MMLS yet remains underexplored.
-Despite the involvement of various modalities, MMLS may exhibit a bias towards relying on a single modality for decision-making Ma et al.
-(2023).
-Recently, few pilot studies have emerged focusing on foundation MMLS.
+We discuss this problem from two perspectives: firstly, the phenomenon of miscalibration in MMLS, and secondly, the strategies for achieving calibration and quantifying uncertainty within these systems.
+5.1. Miscalibration in MMLS
+Miscalibrations in AI systems can manifest as overconfidence, where models inaccurately assign high probabilities to incorrect predictions, or underconfidence, where models fail to recognize the accuracy of their correct predictions with sufficient certainty.
+Similar to uni-modal learning, miscalibration is a prevalent issue in MMLS yet remains underexplored.
+Despite the involvement of various modalities, MMLS may exhibit a bias towards relying on a single modality for decision-making, resulting in the issue of over-confidence (Ma et al., 2023).
+Recently, few pilot studies have emerged focusing on the foundation MMLS.
 For example, concerns about miscalibration have been raised regarding zero-shot VLMs such as CLIP.
-Oh et al.
-(2023) conducts further studies on fine-tuned VLMs, uncovering that traditional fine-tuning strategies can significantly affect model calibration.
+(Oh et al., 2023) conducts further studies on fine-tuned VLMs, uncovering that traditional fine-tuning strategies can significantly affect model calibration.
 Ensuring MMLS are well-calibrated is vital for their safe use, with research concentrating on model calibration and reliable uncertainty estimation (specifically when models are not well-calibrated).
-Calibrating MMLS.
+5.2. Calibrating and Estimating Uncertainty
+A well-calibrated model should make predictions with high uncertainty when its accuracy is in doubt, and conversely, it should predict with confidence when it is likely to be accurate.
 The diversity of modalities in MMLS presents both challenges and potential breakthroughs for the development of new approaches for calibrating MMLS.
-A regularization-based technique CML is proposed for better calibration of MMLS, which calibrates the relationship between confidence and the number of modalities involved in multimodal learning Ma et al.
-(2023).
-Estimating uncertainty in MMLS.
+For instance, a regularization-based technique CML is proposed for better calibration of MMLS, which calibrates the relationship between confidence and the number of modalities involved in multimodal learning (Ma et al., 2023).
 Recent uncertainty estimation research primarily focuses on Bayesian and Conformal Prediction (CP) methods, with extensive work in unimodal settings but limited studies in multimodal contexts.
 Bayesian-based approaches place a distribution over model parameters and employ marginalization for predictive distribution.
-Zhang et al.
-(2023b) pioneers the application of Dempster-Shafer theory to model epistemic uncertainty for each modality and introduces an uncertainty-aware weighting method for dynamic modality fusion.
+There are generally two types of predictive uncertainty: aleatoric uncertainty and epistemic uncertainty.
+(Zhang et al., 2023b) pioneers the application of Dempster-Shafer theory (Shafer, 1976) to model epistemic uncertainty for each modality and introduces an uncertainty-aware weighting method for dynamic modality fusion.
+Meanwhile, Subedar et al.
+(Subedar et al., 2019) introduce a Bayesian framework for audiovisual applications, quantifying modality-wise epistemic uncertainty and deepening Bayesian DNNs by combining deterministic and variational layers.
 While Bayesian-based approaches are widely used, their computational intensity renders them resource-heavy.
-Conversely, CP-based methods have emerged as scalable, efficient, and model-agnostic alternatives for uncertainty estimation.
+Conversely, CP-based methods (Shafer and Vovk, 2008) have emerged as scalable, efficient, and model-agnostic alternatives for uncertainty estimation.
+CP yields statistically valid prediction sets or intervals from any black box model, with the sole assumption of data exchangeability.
 Currently, the application of CP in MMLS is limited.
-Dutta et al.
-(2023) proposes a CP procedure for foundation models using multimodal web data.
+(Dutta et al., 2023) proposes a CP procedure for foundation models using multimodal web data.
 It generates plausibility scores based on modality alignment, and then uses them for Monte Carlo-based CP procedure.
-5.2 Red Teaming MMLS
-The development of foundation models introduces a safety challenge termed ’jailbreak’, where malicious users attempt to circumvent safeguards and instruct the model to produce harmful and unethical responses.
-Consequently, devising strategies to identify and mitigate such phenomena is vital for safety, a primary focus of red teaming efforts Qi et al.
-(2023).
-Historically, research has concentrated on unimodal systems, mostly large language models.
-There has been a recent shift towards exploring these issues within multimodal contexts.
-However, research in this area is still nascent, primarily focusing on VLMs.
-For instance, Qi et al.
-(2023) designs a simple attack that maximizes the generation probability of harmful responses, effectively jailbreaking VLMs and uncovering MMLS’s vulnerabilities.
-More recent efforts, such as Li et al.
-(2024) have begun to establish systematic benchmarks for red teaming in MMLS, paving the way for more in-depth future studies.
-5.3 Discussion
-The field of ensuring MMLS’s reliable output is crucial yet currently underexplored.
-Advancing it requires first addressing MMLS’s common miscalibrations, such as overconfidence or underconfidence, then developing sophisticated calibration and uncertainty quantification methods to improve reliability.
-Besides, red teaming MMLS presents a vital research area.
-Investigating MMLS’s susceptibility to jailbreak attacks and developing countermeasures against such vulnerabilities offers a promising direction for future research.
-6 Interpretability
+5.3. Benchmark
+To measure the reliability of model outputs, a commonly used metric is accuracy vs uncertainty (AvU) (Krishnan and Tickoo, 2020), which considers four types of predictions: accurate and certain, inaccurate and uncertain, accurate and uncertain, inaccurate and certain.
+Other measurements include Expected Calibration Error (ECE) (Guo et al., 2017), Uncertainty Calibration Error (UCE) (Laves et al., 2019), reliability diagrams (Hartmann et al., 2002), and Bayesian active learning by disagreement (BALD) (Houlsby et al., 2011).
+Datasets used for benchmarking can differ widely depending on the task and method.
+5.4. Discussion
+Advancing MMLS’s reliable output stands as a critical yet presently inadequately investigated area of research.
+To improve reliability, it is necessary to address the prevalent issues of miscalibration that these systems often exhibit.
+One of the major challenges in this endeavor involves tackling the intricate interactions between different modalities.
+Although this complexity and variance contribute to miscalibration, it also presents opportunities for effective calibration and precise uncertainty estimation.
+Current multimodal calibration and uncertainty quantification primarily center around Bayesian-based approaches, which tend to be computationally intensive.
+The emerging CP-based method offers a scalable and model-agnostic alternative.
+However, its application in MMLS is still in its infancy, highlighting a ripe area for further investigation.
+6. Controllability -Explainability
 Being able to comprehend and explain the decisions made by the system is critical for the controllability of MMLS.
-In this section, we offer a concise overview of Multimodal Explainable Artificial Intelligence (MXAI) Rodis et al.
-(2023), categorizing MXAI into ante-hoc and post-hoc approaches.
-6.1 Multimodal XAI
-Ante-hoc MXAI.
+The complexity of MMLS makes it difficult to interpret, opaque, and black-box models with little or no understanding of their internal states and decision-making process (Joshi et al., 2021).
+Gaining meaningful knowledge and understanding of how and why the model arrived at a particular decision or outcome is crucial in model explainability, making it one of the important evaluation metrics In this section, we offer a concise overview of Multimodal Explainable Artificial Intelligence, categorizing MXAI into ante-hoc and post-hoc approaches.
+A comprehensive review of MXAI can be found in (Rodis et al., 2023)
+6.1. Multimodal XAI
+6.1.1. Ante-hoc MXAI
 Ante-hoc methods integrate interpretability modules directly into the primary model for simultaneous learning.
 In multimodal contexts, emphasizing the modeling of cross-modal interactions and assessing modalities’ relevance is essential.
-For example, Li et al.
-(2018) introduces a multitask learning approach for VQA that concurrently trains an answer predictor and a textual explanation generator, leveraging attention mechanism for cross-modal dynamics.
+For example, (Li et al., 2018) introduces a multitask learning approach for VQA that concurrently trains an answer predictor and a textual explanation generator, leveraging an attention mechanism for cross-modal dynamics.
 Recently, research on multimodal foundation models has emerged.
-Wang et al.
-(2023) devises an interactive multi-agent VQA framework based on VLMs.
-This framework consists of a seeker and responder that actively generate and exchange information and candidate responses related to the query, and an integrator that synthesizes candidates and generated hypotheses from the seeker and responder to formulate the final answer.
-Post-hoc MXAI.
-Post-hoc methods generally refers to training an interpretability module based on the output of the primary model.
-To address fundamental challenges of capturing interactions across modalities, DIME Lyu et al.
-(2022) proposes to disentangle the black-box model into unimodal contributions and multimodal interactions.
+(Wang et al., 2023b) devises an interactive multi-agent VQA framework based on VLMs.
+This framework consists of a seeker and responder that actively generates and exchanges information and candidate responses related to the query, and an integrator that synthesizes candidates and generated hypotheses from the seeker and responder to formulate the final answer.
+6.1.2. Post-hoc MXAI
+Post-hoc methods generally refer to training an interpretability module based on the output of the primary model.
+To address the fundamental challenges of capturing interactions across modalities, DIME (Lyu et al., 2022) proposes to disentangle the black-box model into unimodal contributions and multimodal interactions.
 This allows for fine-grained analysis of multimodal models while maintaining generality across arbitrary modalities, model architectures, and tasks.
-Another post-hoc method Panousis and Chatzis (2023) focuses on understanding individual neuron functionality in tasks, utilizing vision-text models and unique network layers to achieve high activation sparsity.
+Another post-hoc method (panousis2023discover) focuses on understanding individual neuron functionality in tasks, utilizing VLMs and unique network layers to achieve high activation sparsity.
 It enables the generation of simple textual descriptions, facilitating a direct investigation of the network’s decision process.
-6.2 Benchmark
+6.2. Benchmark
 MXAI datasets are typically organized based on tasks, I/O types, dates, and brief descriptions.
@@
 Evaluation methods can be categorized into three types: application-grounded, involving domain experts for real-world utility; human-grounded, testing understandability with individuals regardless of expertise; and functionally-grounded, using formal criteria to assess explanation quality.
-Detailed descriptions of MXAI benchmarks can be found in Rodis et al.
-(2023).
-6.3 Discussion
+Detailed descriptions of MXAI benchmarks can be found in (Rodis et al., 2023).
+6.3. Discussion
 Formalizing definitions and terminology is crucial for the MXAI field’s development, as current informal practices and inconsistent definitions hinder progress.
@@
 For instance, in contrast to textual explanations, the lack of corresponding ground-truth visual explanation data hampers the development of related explanation methods, especially in supervised learning scenarios.
-7 Fairness
-MMLS incorporates data from various sources like image, text, and audio, presenting unique fairness challenges.
-Prior work indicates that biases in MMLS stem from individual modalities, which could then be amplified during later modality fusion process Yan et al.
-(2020).
+7. Fairness
+MMLS incorporates data from various sources, presenting unique fairness challenges.
+Prior work indicates that biases in MMLS stem from individual modalities, which could then be amplified during later modality fusion process (Yan et al., 2020).
 It is crucial to identify and address the specific fairness challenges and underlying causes in MMLS to effectively mitigate discrimination.
-7.1 Improving Fairness in MMLS
+7.1. Improving Fairness in MMLS
 Addressing fairness in multimodal scenarios is more challenging compared to unimodal due to cross-modality correlation.
-Given the limited research on MMLS fairness, we categorize fairness algorithms based on whether adjustments occur during model training, grouping pre-processing and post-processing together, separate from in-processing techniques.
-7.1.1 Addressing data bias
+Following categorization in uni-modal fair machine learning, we organize fairness algorithms in MMLS into three groups: pre-processing, in-processing, and post-processing.
+7.1.1. Pre-processing
 Pre-processing techniques operate on original input data to remove its inherent bias.
-For instance, in Automated Video Interviews (AVIs) Booth et al.
-(2021), researchers tackle gender bias in verbal, paraverbal, and visual communication.
-They introduces two effective strategies, namely gender-norming and iterative gendered predictor removal, to erase gender-related information from the feature set.
-Post-processing techniques are performed after training by accessing a holdout set which is not involved during model training.
-For instance, Chen et al.
-(2020) determines group-specific random thresholds based on the intersection of group-specific ROC curves.
-By adding this to the clinical MMLS outputs, the prediction bias can be removed.
-Notably, post-processing methods are exclusively usable in black-box settings where the model and training data are inaccessible Although pre-processing and post-processing are more generalizable, easier to implement, and less time-consuming, they may lead to uncertain outcomes if not customized for specific algorithms.
-7.1.2 Mitigating bias during training
+For instance, In the context of Automated Video Interviews, Booth et al.
+(Booth et al., 2021) investigate and tackle gender bias by analyzing verbal, paraverbal, and visual features.
+They present two strategies to mitigate gender-related bias within the feature set: first, gender-norming, which involves separately normalizing the features of men and women before training, and second, iterative gendered predictor removal, a process that systematically eliminates the predictors most indicative of gender information.
+This study also indicates that combining multiple modalities without addressing biases hardly boosts accuracy and tends to amplify unfairness compared to unimodal models.
+Nonetheless, bias mitigation techniques like feature reduction (Schmitz et al., 2022) can achieve an admissible fairness-accuracy trade-off.
+7.1.2. In-processing
 In-processing techniques involve devising training algorithms to mitigate discrimination.
-Compared to above techniques, in-processing methods tend to achieve superior trade-off between fairness and task performance.
-For instance, FairCVtest Pena et al.
-(2020) offers a fictional recruitment testbed demonstrating how biases can be affected by unimodal fators.
+Compared to the pre-processing techniques, in-processing methods tend to achieve superior trade-off between fairness and task performance.
+For instance, FairCVtest (Pena et al., 2020) offers a fictional recruitment testbed demonstrating how biases can be affected by unimodal factors.
 It further ensures fair multimodal outcomes by generating agnostic representations based on SensitiveNets learning strategy.
-However, addressing unfairness in multimodal modeling involves more than just tackling unimodal biases.
-FMMRec Chen et al.
-(2023a) finds that multimodal recommend systems’ bias stems from repetitively identifying sensitive information across modalities.
+However, addressing unfairness in MMLS involves more than just tackling unimodal biases.
+FMMRec (Chen et al., 2023a) finds that multimodal recommend systems’ bias stems from repetitively identifying sensitive information across modalities.
 Adversarial learning method is then employed to disentangle and filter out sensitive information from modality representations.
 Other approaches involve integrating the aforementioned techniques.
-For instance, Yan et al.
-(2020) mitigates gender and race biases through a combination of adversarial learning and data balancing, employing both in-processing and post-processing strategies.
-7.2 Benchmark
-Datasets for fair MMLS are limited due to privacy concerns and the inherent complexity of multimodal data.
-Various fairness notions (e.g., Equalized Odds and Statistical Parity Mehrabi et al.
-(2021)) in MMLS draw inspiration from concepts established in unimodal settings.
-Key fairness categories encompass individual fairness, focusing on individual samples, and group fairness, targeting protected groups such as Black or Female Chen et al.
-(2020).
-In the context of multilingual multimodal models, additional fairness notions like Multilingual Individual and Multilingual Group Fairness have been introduced Wang et al.
-(2021).
-The former necessitates similar semantics across languages, linking text snippets to images, while the latter measures equalized predictive performance across languages.
-7.3 Discussions
+For instance, (Yan et al., 2020) mitigates gender and race biases through a combination of adversarial learning and data balancing, employing both in-processing and post-processing strategies.
+7.1.3. Post-processing
+Post-processing techniques are performed after training by accessing a holdout set not involved during training.
+For instance, (Chen et al., 2020) determines group-specific random thresholds based on the intersection of group-specific ROC curves.
+By adding this to the clinical MMLS outputs, the prediction bias can be removed.
+Notably, post-processing methods are exclusively usable in black-box settings where the model and training data are inaccessible.
+Although pre-processing and post-processing are more generalizable, easier to implement, and less time-consuming, they may lead to uncertain outcomes if not customized for specific algorithms(Mehrabi et al., 2021).
+7.2. Benchmark
+Datasets crucial for ensuring fairness in MMLS are limited due to privacy concerns and the intricate nature of multimodal data.
+Given the scarcity of research in the field, we highlight several representative datasets for reference.
+MovieKens (Ni et al., 2023) is a widely used multimodal benchmark dataset for movie recommendation.
+IEMOCAP (Busso et al., 2008) is a emotion recognition dataset covering audio, text, and video.
+Additionally, the clinical database MIMIC-III (Johnson et al., 2016) and the image-text corpus Multi30K (Elliott et al., 2016) also contribute to fairness research in clinical data and multilingual contexts, respectively.
+Notions of fairness vary in different systems.
+Various fairness notions (e.g., Equalized Odds and Statistical Parity (Mehrabi et al., 2021)) in MMLS draw inspiration from concepts established in unimodal settings.
+Key fairness categories encompass individual fairness, focusing on individual samples, and group fairness, targeting protected groups such as Black or Female (Chen et al., 2020).
+In the context of multilingual MMLS, additional fairness metrics like Multilingual Individual and Multilingual Group Fairness have been introduced (Wang et al., 2021).
+The former necessitates similar semantics across languages, linking text to images, while the latter measures equalized predictive performance across languages.
+7.3. Discussions
 In contrast to the abundance of datasets available for fairness studies in unimodal scenarios, there is a notable absence of reliable datasets for fairness studies in MMLS.
-This gap is attributed to the intricate nature of modal information, privacy concerns, and the diverse system considerations related to fairness.
-Existing work, such as Schmitz et al.
-(2022), compares biases introduced by different modalities and their combinations, yet falls short of delving into the fundamental causes of multimodal unfairness and proposing effective solutions.
+This gap is attributed to the intricate nature of multi-modal information, privacy concerns, and the diverse system considerations related to fairness.
+Existing work, such as (Schmitz et al., 2022), compares biases introduced by different modalities and their combinations, yet falls short of delving into the fundamental causes of multimodal unfairness and proposing effective solutions.
+One may explore the causes from an uncertainty perspective (Tahir et al., 2023).
 Furthermore, there is a crucial need for the introduction of new definitions of fairness.
-While some notable work has suggested some Wang et al.
-(2021), it lacks a thorough explanation of why this new definition is necessary or when it should be applied.
-8 Open Challenges
-Our review highlights several key challenges in MMLS safety as briefly illustrated below. We aim to offer insights that can serve as valuable guidance for future research.
+While some notable work has suggested some (Wang et al., 2021), it lacks a thorough explanation of why this new definition is necessary or when it should be applied.
+8. Privacy
+8.1. Privacy Issues in MMLS
+Significant privacy concerns regarding MMLS have emerged due to their high risk of memorizing and leaking private information (Chen et al., 2023b; Naseh et al., 2023).
+Membership Inference Attacks (Ko et al., 2023), for example, could reveal an individual’s health status by checking if their record is in a multimodal disease database.
+Besides, reconstruction attacks may exploit MMLS’s memorized information to extract private data, such as identifying residents in a geospatial system using modality alignment information between street images and resident details (Rao et al., 2023).
+The recent emerging large MMLS following the pretrain-then-finetune paradigm further intensifies privacy concerns.
+Each stage of these models, pre-training, fine-tuning, and prompt interaction, carries a high risk of privacy leakage (Rao et al., 2023).
+8.2. Preserving Privacy in MMLS
+Relevant research primarily focuses on differential privacy, machine unlearning, and federated learning.
+8.2.1. Differentially private training.
+The differential privacy (DP) technique, recognized as a gold standard in privacy preservation, ensures data protection by introducing randomness into its computational processes.
+DP in MMLS remains understudied.
+Multimodal DP was first introduced to address the heightened risk of privacy leakage arising from MMLS memorizing sensitive training data information (Huang et al., 2023).
+In this study, DP-CLIP is proposed to safeguard vision-language tasks by learning differentially private image and text representations.
+It employs per-batch, instead of per-sample, noisy stochastic gradient descent to train the encoder, considering that the CLIP loss requires contrasting data from different pairs.
+8.2.2. Machine unlearning.
+Machine unlearning aims at selectively erasing information (e.g., sensitive information) from a trained model without the need for complete retraining.
+Progress in machine unlearning has advanced in unimodal learning, but its application in MMLS is still in its infancy.
+Compared to unimodal scenarios, the key challenge lies in effectively addressing the inherent interdependencies across modalities.
+One work (Shaik et al., 2023) presents a modality-specific attention mechanism that generates individual attention scores per modality, averages them across modalities for each data point to determine its final score, and then decides on data removal and model parameter updates based on this score.
+Recently a more comprehensive framework for multimodal unlearning has been proposed (Cheng and Amiri, 2023), encompassing modality decoupling, unimodal knowledge retention, and multimodal knowledge retention.
+Its first module ensures that individual data modality in deleted pairs are treated as unrelated by the model, while the latter two focus on preserving essential knowledge to maintain the model performance.
+8.2.3. Federated learning.
+Federated learning enables collaborative model training without the need to share clients’ local data, offering a more privacy-enhancing approach compared to centralized training methods.
+The MMFed framework, proposed in (Xiong et al., 2022), addresses modality discrepancies in MMLS settings through a unified multimodal federated learning approach.
+It utilizes a cross-attention mechanism to identify key correlations across modalities on local clients, then uploads the parameters of both the classifier and attention module to a central server.
+CreamFL further advances multimodal federated learning by enabling larger server models and aggregation of clients with diverse architectures and data modalities, thereby aligning more closely with practical scenarios (Yu et al., 2023).
+8.3. Benchmark
+Current memorization measurements lack standardization and are primarily restricted to VLMs.
+Common metrics include Frechet Inception Distance score (Somepalli et al., 2023) and Memorization score (Zhang et al., 2023a).
+Machine unlearning evaluation parallels that in unimodal settings, including attack-based methods like Membership Inference Robustness, and accuracy-based metrics like Mean Squared Error, alongside others like Activation Distance and Retraining Time (Nguyen et al., 2022).
+Benchmarking multimodal federated learning is still in its early stages.
+FedMultimodal is the first systematic benchmark (Feng et al., 2023), offering 10 representative datasets from 5 key MMLS applications (emotion recognition, action recognition, human activity recognition, healthcare, and social media) and providing a framework covering the pipeline of multimodal federated learning for effective assessment.
+8.4. Discussion
+Due to the diversity and richness of input data modalities, the potential risk of privacy leakage is concerning in MMLS.
+This involves issues such as the recovery of removed sensitive information after multimodal fusions.
+Besides, the increasing scale of MMLS further exacerbates the challenge of memorization.
+Despite these concerns, privacy issues and their underlying causes are still not thoroughly explored, lacking formal conceptualizations and investigations.
+One potential direction is to extend findings in (Zhao et al., 2023) to multimodal data.
+9. Open Challenges
+We highlights several challenges in MMLS safety, aiming to offer insights that can serve as valuable guidance for future research.
 Memorization Problem in MMLS.
@@
 Safety Challenges in Multimodal LLMs.
-Multimodal LLMs (e.g., GPT 4V) are designed to understand and generate content across different modalities.
+Multimodal LLMs (e.g., GPT 4V (Achiam et al., 2023)) are designed to understand and generate content across different modalities.
 Safeguarding multimodal LLMs is very challenging due to various reasons including the additional safety rift induced by different data modalities, often black-box API access, unknown data sources, and emerging LLM safety issues such as hallucination, data leakage, and jailbreak.